# LoaQ: Layer-wise Output Approximation Quantization

### v1

@@ 3.2 Sub-layer Output Approximation
Figure 3: Results under 3-bit quantization on LLaMA 3.2 1B, showing the mean squared error (MSE) of sub-layers’ outputs without outlier tokens. The left plot shows the MSE of unnormalized outputs, and the right plot shows the MSE of normalized outputs.
3.3 Normalized Sub-layer Output Approximation
In this subsection, we shift the focus from approximating the blue circle 3 to approximating the orange circle 4 in Figure 1, aiming to make the post-quantization results closely match their pre-quantization counterparts by approximating the normalized outputs of the corresponding sub-layer.
Based on the above understanding of the Transformer architecture, the presence of residual connections leads to the accumulation of quantization error as the depth increases, as shown in the left plot of Figure 3. This raises the question of how the model remains stable despite the accumulation of quantization error. Upon closer examination, we observed that an RMSNorm layer is consistently applied before the input to each sub-layer module and the final language model head. As illustrated in the right plot of Figure 3, this normalization mechanism prevents the mean squared error (MSE) between normalized outputs from growing beyond a certain depth.
@@ 3.3 Normalized Sub-layer Output Approximation
Since a normalization layer follows each sub-layer, it is the normalized output rather than the raw sub-layer output that directly influences subsequent linear projections. As such, aligning the normalized outputs of the quantized and original models becomes critical for maintaining the model’s functional integrity. To address this, we refine the objective to explicitly target the normalized sub-layer outputs:
$\arg\min_{Q}\|\operatorname{Norm}(\hat{h}+\hat{X}Q)-\operatorname{Norm}(h+XW)\|_{F}^{2}$ (11)
To simplify the expression, we write the normalization term as $\operatorname{Norm}(\hat{h}+\hat{X}Q)=R(\hat{h}+\hat{X}Q)(\hat{h}+\hat{X}Q)$, where both the scaling function $R(\cdot)$ and the input contain the quantized parameter $Q$. To reduce the number of unknowns and decouple the quantized terms, we replace $Q$ with the original weight $W$ inside the scaling function $R(\cdot)$, and can then be transformed into the form:
$\arg$ $\min_{Q}\|\big{(}R(\hat{h}+\hat{X}W)\hat{h}+(R(\hat{h}+\hat{X}W)\hat{X})Q\big{)}$
$-\big{(}R(h+XW)h+(R(h+XW)X)W\big{)}\|_{F}^{2}$ (12)
This yields the same form as objective (9), allowing us to apply the same solution.
@@ 4.1 Baselines and Evaluation
We experiment with our proposed LoaQ on three commonly-used LLM families, covering different sizes: LLaMA 2 (7B, 13B, 70B) (Touvron et al. 2023), LLaMA 3 (8B, 70B) (Grattafiori et al. 2024) and Qwen 3 (8B, 14B, 32B) (Yang et al. 2025). Under the weight-only quantization setting, we compare LoaQ with the baseline GPTQ (Frantar et al. 2023) and two error compensation methods, Qronos (Zhang et al. 2025) and GPTAQ (Li et al. 2025). We also evaluate scenarios that incorporate effective quantization techniques such as NeUQI (Lin, Hu, and Wan 2025) and the Hadamard transform, in order to test whether LoaQ remains effective under these enhancements. In addition, LoaQ is assessed under the weight-activation quantization setting. A brief introduction to these techniques is provided in Appendix B. Following the previous work, all the quantized models are evaluated by measuring perplexity on the WikiText2 (Wiki2(Merity et al. 2017) and C4 (Raffel et al. 2020) validation sets, and zero-shot accuracy on five benchmarks: ARC-easy (ArcE), ARC-challenge (ArcC(Clark et al. 2018), PiQA (Bisk et al. 2020), HellaSwag (HellaS(Zellers et al. 2019) , and WinoGrande (WinoG(Sakaguchi et al. 2021).
@@ 4.2 Implementation Detail
For weight-only quantization, we evaluate the 2-bit setting and the 3-bit setting, both using channel-wise quantization for weights. For weight-activation quantization, we adopt the W4A4 setting, where weights are quantized to 4 bits channel-wise and activations to 4 bits token-wise dynamically, with quantization parameters determined by the Min-Max method (Jacob et al. 2017). During calibration, we draw 128 samples of 2048 tokens each from the C4 dataset for LLaMA 2 and 3, and Qwen 3 families, whereas other studies may use WikiText2. We perform a hyperparameter grid search on LLaMA 3.2 1B and Qwen 3 1.7B, where $\alpha$ ranges from 0 to 1 in increments of 0.1 and $\beta$ ranges from 0 to 1 in increments of 0.05, and transfer the optimized configurations to the larger models. Additional tuning on larger models is conducted only when the transferred settings perform poorly. All experiments are conducted on a single NVIDIA A40 GPU. More implementation details are provided in Appendix C.
@@ 4.3 Main Result
As the bit-width increases, the performance gap between methods narrows, as shown in Table 1 of Appendix D. LoaQ continues to show clear advantages under the 3-bit setting, demonstrating robustness across model architectures, while in some cases performing on par with others and still offering slight benefits in specific scenarios.
@@ 4.4 Discussion
Ablation Study
We further evaluate the individual contributions of the Sub-layer Output Approximation (SOA) and the Normalized Sub-layer Output Approximation (NOA) in LoaQ through an ablation study on LLaMA 2 7B under 2-bit channel-wise quantization. We examine three configurations: removing SOA, removing NOA, and removing both SOA and NOA. The results in Table 5 demonstrate that both SOA and NOA provide clear benefits. Moreover, as shown in Figure 4, the top two plots without the Hadamard transform reveal that SOA markedly reduces the mean squared error of hidden states across all layers, while NOA substantially lowers errors in the early layers. The corrective effect of NOA appears to diminish with increasing layer depth, likely because accumulated quantization errors weaken its impact. In contrast, the bottom two plots with the Hadamard transform show a considerable overall improvement. Under this setting, the influence of NOA is less directly reflected in the error, while SOA still exhibits a clear reduction in errors. These results confirm that both SOA and NOA are effective, and their combination, especially when paired with the Hadamard transform, further enhances quantization performance.
@@
Table 5: Ablation study of the SOA and NOA modules with and without the Hadamard Transform on LLaMA 2 7B using 2-bit channel-wise quantization.
Figure 4: Results of 2-bit channel-wise quantization on LLaMA 2 7B, presenting the mean squared error (MSE) of Transformer layer outputs excluding outlier tokens. The top two plots show the MSE without the Hadamard transform, while the bottom two plots show the MSE with the Hadamard transform. For each row, the left plot corresponds to unnormalized outputs, and the right plot corresponds to normalized outputs.
Quantization Overhead
@@ 5 Conclusion
Appendix A Proofs
A.1 General Form
Lemma 1.
If $X^{\top}X$ is invertible, then
$\arg\min_{Q}\|XQ-Y\|_{F}^{2}\;=\;\arg\min_{Q}\Bigl{\|}X\bigl{(}Q-Q^{*}\bigr{)}\Bigr{\|}_{F}^{2},$ (13)
where
$Q^{*}=(X^{\top}X)^{-1}X^{\top}Y.$ (14)
Proof.
Consider the functions
$f(Q)$ $=\|XQ-Y\|_{F}^{2},$ (15)
$g(Q)$ $=\|X(Q-Q^{*})\|_{F}^{2}.$ (16)
We will demonstrate that $f$ and $g$ have identical gradients with respect to $Q$, and hence share the same minimizer.
First, expand $f(Q)$ in trace form:
$f(Q)$ $=\mathrm{tr}\bigl{(}(XQ-Y)(XQ-Y)^{\top}\bigr{)}$ (17)
$=\mathrm{tr}(Q^{\top}X^{\top}XQ)-2\,\mathrm{tr}(Q^{\top}X^{\top}Y)+\mathrm{tr}(Y^{\top}Y).$ (18)
Differentiating with respect to $Q$ gives
$\nabla_{Q}f(Q)=2X^{\top}XQ-2X^{\top}Y=2X^{\top}X(Q-Q^{*}).$ (19)
Next, express $g(Q)$ in trace form:
$g(Q)$ $=\mathrm{tr}((Q-Q^{*})^{\top}X^{\top}X(Q-Q^{*})).$ (20)
Its derivative is
$\nabla_{Q}g(Q)=2X^{\top}X(Q-Q^{*}).$ (21)
Since $\nabla_{Q}f(Q)=\nabla_{Q}g(Q)$ for all $Q$, the functions $f$ and $g$ differ only by a constant and consequently have the same minimizer. Therefore,
$\arg\min_{Q}f(Q)\;=\;\arg\min_{Q}g(Q).$ (22)
A.2 Specific Form
In the methodology section, we aim to prove the following results, where we omit the subscript $Q$ in $\arg\min_{Q}$:
$\arg\min\|\hat{X}Q-XW\|_{F}^{2}$ (23)
$=\arg\min\|\hat{X}\big{(}Q-(\hat{X}^{T}\hat{X})^{-1}(\hat{X}^{T}X)W\big{)}\|_{F}^{2}$ (24)
$=\arg\min\|\hat{X}\big{(}Q-H^{-1}(H+C)W\big{)}\|_{F}^{2}$ (25)
$=\arg\min\|\hat{X}\big{(}Q-(I+H^{-1}C)W\big{)}\|_{F}^{2}$ (26)
$\arg\min\|(\hat{h}+\hat{X}Q)-(h+XW)\|_{F}^{2}$ (27)
$=\arg\min\|\hat{X}Q-(XW+h-\hat{h})\|_{F}^{2}$ (28)
$=\arg\min\|\hat{X}\big{(}Q-((\hat{X}^{T}\hat{X})^{-1}(\hat{X}^{T}X)W$
$\qquad\qquad\qquad\qquad+(\hat{X}^{T}\hat{X})^{-1}\hat{X}^{T}(h-\hat{h}))\big{)}\|_{F}^{2}$ (29)
$=\arg\min\|\hat{X}\big{(}Q-((I+H^{-1}C)W$
$\qquad\qquad\qquad\qquad\quad+H^{-1}\hat{X}^{T}(h-\hat{h}))\big{)}\|_{F}^{2}$ (30)
Appendix B Quantization Techniques
@@ B.1 GPTQ
Follow LDLQ proposed by QuIP, we revisit GPTQ and propose a more effective derivation approach, which can be formally expressed as follows.
$\|\hat{X}(Q-W)\|_{F}^{2}$ $=\operatorname{tr}\big{(}(Q-W)^{\top}(\hat{X}^{\top}\hat{X})(Q-W)\big{)}$
$=\operatorname{tr}\big{(}(Q-W)^{\top}H(Q-W)\big{)}$
$=\operatorname{tr}\big{(}(Q-W)^{\top}L^{\top}L(Q-W)\big{)}$
$=\|L(Q-W)\|_{F}^{2},$ (31)
where $L$ is the Cholesky factor of of $H$ and is a lower triangular matrix. The term $\|L(Q-W)\|_{F}^{2}$ can be decomposed into the sum of squared errors of each row. Notably, the $i$-th row of L(QW)L(Q-W) is a linear combination of the first $i$ rows of QWQ-W. This property enables sequential row-wise quantization. Consequently, we derive
$Q$ $=\operatorname{Quant}\big{(}W-D^{-1}(L-D)(Q-W)\big{)},$ (32)
where $D$ denotes the diagonal matrix formed by the diagonal elements of $L$, and $\operatorname{Quant}(\cdot)$ maps each item to the nearest value in the quantization grid.
To ensure that $H$ is positive definite, a sufficiently large term $\lambda I$ is typically added. This is essentially equivalent to solving
$\arg\min_{Q}\|\hat{X}(Q-W)\|_{F}^{2}+\lambda\|Q-W\|_{F}^{2}.$ (33)
Additionally, we follows the GPTQ heuristic of reordering each row based on the diagonal values of $H$ in descending order.
B.2 Hadamard transform
The Hadamard transform [^1] See https://en.wikipedia.org/wiki/Hadamard˙matrix, which is used in methods such as QuaRot and QuIP#, smooths the distributions of activations and weights, enabling more effective quantization while maintaining high efficiency due to the special structure of the Hadamard matrix, which allows fast computation. Specifically, it employs an orthogonal matrix $R$ with entries $\pm 1$, enabling the transformations $X^{\prime}=XR$ and $W^{\prime}=R^{-1}W$. Here, $X^{\prime}$ and $W^{\prime}$ serve as the transformed activations and weights that are subsequently used for quantization. When the transform is applied to a vector of length $n$, the computational complexity is $O(n\log n)$ on a single worker and can be reduced to $O(\log n)$ with sufficient workers.
B.3 NeUQI
Appendix C Implementation Details
C.1 LoaQ Pipeline
The following algorithm describes the overall LoaQ pipeline.
Algorithm 1 LoaQ Pipeline
1:Weights $\{W_{\mathrm{in}}^{(\ell)},W_{\mathrm{out}}^{(\ell)}\}_{\ell=1}^{L}$, calibration inputs $\{I^{(i)}\}_{i=1}^{S}$, bit‐width $k$, adjustment parameters $\alpha,\beta\in[0,1]$
2:Quantized weights $\{Q^{(\ell)}\}_{\ell=1}^{L}$
3:Run forward to collect embedding outputs $h,\hat{h}$
4:for $\ell=1$ to $L$ do
5:  for $W$ in $\{W_{\mathrm{in}}^{(\ell)},W_{\mathrm{out}}^{(\ell)}\}$ do
6:   Run forward to collect activation $X,\hat{X}$
8:     $H\leftarrow\hat{X}^{\top}\hat{X}$
[…]
10:     $\Delta=\textbf{0}$
[…]
12:     $s\leftarrow R(h+XW)$
13:     $\hat{s}\leftarrow R(h+\hat{X}W)$ $\triangleright$ Rescale for each token
14:     $X^{\prime}\leftarrow sX$
15:     $\hat{X}^{\prime}\leftarrow\hat{s}\hat{X}$
16:     $H\leftarrow\hat{X}^{\prime\top}\hat{X}^{\prime}$
17:     $C\leftarrow\hat{X}^{\prime\top}(X^{\prime}-\hat{X}^{\prime})$
18:     $\Delta\leftarrow H^{-1}\hat{X}^{\prime\top}\bigl{(}sh-\hat{s}\hat{h}\bigr{)}$
19:   end if
20:   $\widetilde{W}\leftarrow\bigl{(}I+\alpha H^{-1}C\bigr{)}W+\beta\Delta$
21:   Quantize: $Q\leftarrow\mathrm{QuantMethod}(\widetilde{W};H;k)$
22:  end for
23:  Run forward to get updated sub-layer outputs $h,\hat{h}$
24:end for
25:return $\{Q^{(\ell)}\}_{\ell=1}^{L}$
C.2 Hyperparameter
We perform a hyperparameter grid search by first tuning $\alpha$ and then tuning $\beta$. Specifically, $\alpha$ ranges from 0 to 1 in increments of 0.1 and $\beta$ ranges from 0 to 1 in increments of 0.05.
@@ C.2 Hyperparameter
$\beta$ { $\in\{\,0.05\times j\mid j=0,1,2,\ldots,20\,\}.$ (35)
Appendix D Supplementary Results

### v2

@@ 4.4 Tuning Parameters
12:     $s^{\prime}\leftarrow R(h^{\prime}+X^{\prime}W)$
13:     $s\leftarrow R(h+XW)$ $\triangleright$ Rescale for each token from Section 4.3
14:     $H\leftarrow(sX)^{\top}(sX)$
15:     $C\leftarrow(sX)^{\top}(s^{\prime}X^{\prime}-sX)$
16:     $\Delta\leftarrow H^{-1}(sX)^{\top}\bigl(s^{\prime}h^{\prime}-sh\bigr)$ $\triangleright$ Correction term from Section 4.2
17:   end if
18:   $\widetilde{W}\leftarrow\bigl(I+\alpha H^{-1}C\bigr)W+\beta\Delta$
19:   Quantize: $Q^{(\ell)}\leftarrow\mathrm{GPTQ}(\widetilde{W};H;k)$
20:  end for
21:  Run forward to get updated sub-block outputs $h,h^{\prime}$
22:end for
23:return $\{Q^{(\ell)}\}_{\ell=1}^{L}$
@@ 5.1 Baselines and Evaluation
We experiment with our proposed LoaQ on three commonly-used LLM families, covering different sizes: LLaMA 2 (7B, 13B, 70B) Touvron et al. (2023), LLaMA 3 (8B, 70B) Grattafiori et al. (2024) and Qwen 3 (8B, 14B, 32B) Yang et al. (2025). Under the weight-only quantization setting, we compare LoaQ with the baseline GPTQ Frantar et al. (2023) and two error compensation methods, Qronos Zhang et al. (2025) and GPTAQ Li et al. (2025). We also evaluate scenarios that incorporate effective quantization techniques such as NeUQI Lin et al. (2025) and the Hadamard transform to assess the effectiveness of LoaQ under these enhancements. In addition, LoaQ is assessed under the weight-activation quantization setting. A brief introduction to these techniques is provided in Appendix A. Following the previous work, all the quantized models are evaluated by measuring perplexity on the WikiText2 (Wiki2Merity et al. (2017) and C4 Raffel et al. (2020) validation sets, and zero-shot accuracy on five benchmarks: ARC-easy (ArcE), ARC-challenge (ArcCClark et al. (2018), PiQA Bisk et al. (2020), HellaSwag (HellaSZellers et al. (2019) , and WinoGrande (WinoGSakaguchi et al. (2021).
@@ 5.2 Implementation Detail
For weight-only quantization, we evaluate the 2-bit channel-wise quantization and the 3-bit channel-wise quantization. For weight-activation quantization, we adopt the W4A4 setting, where weights are quantized to 4 bits channel-wise and activations to 4 bits token-wise dynamically, with quantization parameters determined by Min-Max Jacob et al. (2017). During calibration, following prior work, we draw 128 samples of 2048 tokens each from the C4 dataset for the LLaMA 2, LLaMA 3, and Qwen 3 families, with a fixed random seed to ensure reproducibility. We use the LM Evaluation Harness (Gao et al., 2024) to evaluate zero-shot accuracy on five benchmarks. All experiments are conducted on NVIDIA A40. Our proposed LoaQ reformulates the problem as Eq. 3.1. We use GPTQ in our experiments, but the formulation also supports other optimization methods. More implementation details are provided in Appendix B.
@@ 5.3 Main Result
As the bit-width increases, the performance gap between methods narrows, as shown in Table 7 in Appendix. LoaQ continues to show clear advantages under the 3-bit setting, demonstrating robustness across model architectures, while in some cases performing on par with others and still offering slight benefits in specific scenarios.
@@ 5.4.2 Parameter Initialization
5.4.3 Weight-Activation Quantization
5.4.4 Ablation Study
We further evaluate the individual contributions of the Sub-block Output Approximation (SOA) and the Normalized Sub-block Output Approximation (NOA) in LoaQ through an ablation study on LLaMA 2 7B under 2-bit channel-wise quantization. We examine three configurations: removing SOA, removing NOA, and removing both SOA and NOA. The results in Table 5 demonstrate that both SOA and NOA provide clear benefits. Moreover, as shown in Figure 4, the top two plots without the Hadamard transform reveal that SOA markedly reduces the mean squared error of hidden states across all layers, while NOA substantially lowers errors in the early layers. The corrective effect of NOA appears to diminish with increasing layer depth, likely because accumulated quantization errors weaken its impact. In contrast, the bottom two plots with the Hadamard transform show a considerable overall improvement. Under this setting, the influence of NOA is less directly reflected in the error, while SOA still exhibits a clear reduction in errors. These results confirm that both SOA and NOA are effective, and their combination, especially when paired with the Hadamard transform, further enhances quantization performance.
@@ 5.4.4 Ablation Study
GPTQ 6875 2237 35.51 759 277 34.95
Table 5: Ablation study of the SOA and NOA with and without the Hadamard Transform on LLaMA 2 7B using 2-bit channel-wise quantization.
Figure 4: Results of 2-bit channel-wise quantization on LLaMA 2 7B, presenting the mean squared error (MSE) of Transformer block outputs excluding outlier tokens. The top two plots show the MSE without the Hadamard transform, while the bottom two plots show the MSE with the Hadamard transform. For each row, the left plot corresponds to unnormalized outputs, and the right plot corresponds to normalized outputs.
5.4.5 Quantization Overhead
@@ 6 Conclusion
Limitations
In this work, we focus on corrections within the layer-wise PTQ framework, particularly GPTQ. Accordingly, the evaluation is restricted to the layer-wise PTQ framework, and the resulting conclusions may not generalize to methods outside the layer-wise PTQ setting. Moreover, the LoaQ method alone still exhibits a performance gap compared to full-precision (FP16/BF16) models and therefore requires other complementary techniques for support. In addition, the performance of quantized models can vary significantly depending on the distribution of the training data, and their behavior in more specific application scenarios remains unclear.
Ethical considerations
This work does not raise any ethical concerns. All open-source large language models and datasets used in our work are publicly available. All resources are used in compliance with their respective licenses and solely for research purposes. We carefully review the datasets used in our work to ensure that they do not contain personally identifiable information or offensive content. For the use of AI assistants, we employ them to polish descriptive content and generate plotting code.
Appendix A Quantization Techniques
@@
A.1 Qronos
The Qronos algorithm Zhang et al. (2025) optimizes the same loss function as GPTAQ, as defined in Eq. 4. The formulation of Qronos is
[…]
$\quad\quad+I_{:,2:}(H_{2:,2:})^{-1}C_{2:,:}W,$
$Q$ $=\mathcal{Q}(\widetilde{W}+e^{(1)}(H_{1,1}^{-1}C_{1,:}W)),$ (15)
where $I$ denotes the identity matrix and $e^{(1)}$ denotes the first standard basis vector. Consequently, the term $e^{(1)}H_{1,1}^{-1}C_{1,:}W$ applies a correction to the first row of $\widetilde{W}$ by adding $H_{1,1}^{-1}C_{1,:}W$. In addition, directly applying the formulation in Eq. A.1 yields worse performance than GPTQ. Therefore, Zhang et al. (2025) set $X^{\prime}=X$ at the beginning of each block.
A.2 Hadamard Transform
The Hadamard transform [^2] See https://en.wikipedia.org/wiki/Hadamard_matrix, which is used in methods such as QuaRot and QuIP#, smooths the distributions of activations and weights, enabling more effective quantization while maintaining high efficiency due to the special structure of the Hadamard matrix, which allows fast computation. Specifically, it employs an orthogonal matrix $R$ with entries $\pm 1$, enabling the transformations $X^{\prime}=XR$ and $W^{\prime}=R^{-1}W$. Here, $X^{\prime}$ and $W^{\prime}$ serve as the transformed activations and weights that are subsequently used for quantization. When the transform is applied to a vector of length $n$, the computational complexity is $O(n\log n)$ on a single worker and can be reduced to $O(\log n)$ with sufficient workers.
A.3 NeUQI
Appendix B Implementation Details
@@ B.1 GPTQ
Our proposed LoaQ reformulates the problem as Eq. 1, which we solve using the LDLQ algorithm (Chee et al., 2023), which is mathematically equivalent to GPTQ. To improve numerical stability, we add a small diagonal damping term equal to one percent of the average value of the diagonal entries of $H$ to the Hessian matrix $H$ to ensure that it is invertible. To further improve performance in the 2-bit setting, following Frantar et al. (2023), we execute the LDLQ algorithm by permuting the row order of the weight matrix in descending order of the diagonal entries of $H$, rather than using the original row order.
B.2 Hyperparameters
We perform a hyperparameter grid search by first tuning $\alpha$ and then tuning $\beta$. Specifically, $\alpha$ ranges from 0 to 1 in increments of 0.1 and $\beta$ ranges from 0 to 0.5 in increments of 0.05.
@@ B.2 Hyperparameters
$\beta$ { $\in\{\,0.05\times j\mid j=0,1,2,\ldots,10\,\}.$ (17)
$\alpha=0.4$ and $\beta=0.15$ are empirically selected as effective values based on a search conducted on the LLaMA and Qwen families.
Appendix C Supplementary Results